OpenStudy (anonymous):

sinA (1+tan A)+cos A(1+cot A)=secA+cosec A.prove it!!

OpenStudy (anonymous):

i'd rewrite all the tan, sec, cosecs to sin and cos

OpenStudy (anonymous):

tan = sin/cos cot = cos/sin sec = 1/cos cosec = 1/sin and now try Left Hand Side and Right Hand Side seperately you will see they both are same

OpenStudy (anonymous):

I was going to help her

OpenStudy (anonymous):

if you are still struck, ask ..good luck shruti

OpenStudy (anonymous):

Is she still here anymore?

OpenStudy (anonymous):

thx u both!! :)

OpenStudy (anonymous):

Shruti....you're back :)

OpenStudy (anonymous):

Do you need help with anything else?

OpenStudy (anonymous):

Another identity maybe?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

Okay, hold on a minute...I will post my online classroom link...

OpenStudy (anonymous):

You can join me in my classroom and we can go over a problem.

OpenStudy (anonymous):

I posted my online classroom link above.

OpenStudy (anonymous):

^1+sinA/1-sinA-secA=secA-^1-sinA/1+sinA.

OpenStudy (anonymous):

You can re-post the problem in the classroom and we can go over it.

OpenStudy (anonymous):

Shruti, are you still here?